OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -4 - 2t. Find the instantaneous velocity at t = 6 by finding the derivative.

OpenStudy (anonymous):

(f(a+h) - f(a)) / h

OpenStudy (anonymous):

the velocity is ds/dt. So, differentiate s(t) to get v(t) and then plug the value of t in the expression.

OpenStudy (anonymous):

how do i differentiate?

OpenStudy (anonymous):

\[\frac{ f(a+h) -f(a) }{ h}\]

OpenStudy (anonymous):

I don't think that makes any sense, which one would be f,a and h?

OpenStudy (anonymous):

when you differentiate the constant term will vanish. Any term like \[cx^n\] will become \[ncx^{n-1}\]. Thus 2t will become \[1.2.t^0 = 2.1 = 2\]

OpenStudy (anonymous):

So once we differntiate, I just add/subtract and that'll be my answer?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

rkopen has given you the shortcut to solving this problem. Do you understand

OpenStudy (anonymous):

-4-2 = -6 So -6 is my derivative, correct?

OpenStudy (anonymous):

No -4 term will become 0 since 4 is a constant.

OpenStudy (anonymous):

0-2=-2?